What’s the normal frequency at which DNA-Pol δ/ε introduces mutations in the DNA during replication?

A) 10-8

B) 10-5

C) 10-3
Answer #1

The normal frequency at which DNA Pol δ/ε introduces mutations in DNA during replication is A) 10-8 .

Because in DNA replication newly double stranded DNA is synthesized by DNA polymerase, it contains several subunits, lagging strand is synthesized by DNA Pol δ and leading strand is synthesized by DNA Pol ε .DNA polymerase has high accuracy power (high fidelity), it has 3'--5' exonuclease activity by which DNA Pol can proofread the mismatched base, error sometimes occur about 1 out of every 100,000,000 times = 1/108 = 10-8 times, a mistake occurs during copying which can leads to mutation in DNA.